OpenStudy (anonymous):

find f ' (x) and f '(c) f(x)= (1/3)(2x^3 - 4) value of c = 0

OpenStudy (anonymous):

i'm not really sure how to get started

OpenStudy (across):

\[ f(x)=\frac13\left(2x^3-4\right)\\ f'(x)=2x^2\\ f'(0)=0. \]

OpenStudy (anonymous):

how did you get those primes?

OpenStudy (across):

I used both\[ f(x)=ax^{b}\Longrightarrow f'(x)=abx^{b-1}, \]and\[ f(x)=c\Longrightarrow f'(x)=0. \]

OpenStudy (anonymous):

ah ok, thanks